Circular No. 6008

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994T IN ANONYMOUS GALAXY Corrigendum. P. Challis, Center for Astrophysics, notes that the po-

sition of the supernova reported on IAUC 6007 was calculated incorrectly; the correct position is α = 13^{h}18^{m}56^{s}.16, δ = −1^{o}53′15′′.0 (equinox 1950.0), and the galaxy’s center has end figures 55^{s}.95, 02′′.2.

VERY LARGE ARRAY SKY SURVEY D. Finley, National Radio Astronomy Observatory (NRAO), writes:

“The first radio maps of the sky produced by the 1.4-GHz NRAO Very Large Array Sky Survey (NVSS) now are available to the astronomical community. Observations for the NVSS began in 1993 September and are planned to cover the sky north of δ = −40^{o} (82 percent of the celestial sphere) before the end of 1996. The data products of the NVSS are being released as soon as they are produced and verified. All large NVSS maps that have been verified are available by anonymous FTP, from the host ‘gibbon.cv.nrao.edu’. The NVSS files are in the directory ‘pub/nvss’. Users should first obtain the README file for a description of the data products. The information also may be accessed via WWW. The NRAO master home page URL is ‘http://info.aoc.nrao.edu’. If you have any questions or comments, please contact J. Condon by email at Internet address ‘jcondon@nrao.edu’ or by telephone at 804-296-0322.”

NOVA OPHIUCHI 1994 A. C. Gilmore and P. M. Kilmartin, Mount John University Obser-

vatory, report additional preliminary photometry (cf. IAUC 6002): June 5.657 UT, V = 8.73, U − B = −0.29, B − V = +0.46, V − R_{c} = +0.76, V − I_{c} = +1.28; 6.486, 8.81, −0.30, +0.47, +0.82, +1.30.

Visual magnitude estimates from L. Szentaskó, Veresegyház, Hungary: June 9.006 UT, 9.0; 13.91, 9.7.

PERIODIC COMET TEMPEL 1 (1993c) Total visual magnitude estimates (cf. IAUC 5950): May 2.10 UT,

9.8 (J. E. Bortle, Stormville, NY, 0.32-m reflector); June 5.21, 9.4 (A. Hale, near Orogrande, NM, 10×50 binoculars); 8.88, 9.4 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector).

1994 June 14 Daniel W. E. Green